Be sure to answer all parts. A coal-burning power plant generates 5.5 x 1016 J per year. If the coal yields 38 kJ/g, how many grams of coal was burned each year? x 10 g Enter your answer in scientific notation.
Answer #1

1 year = 5.5*10^16 joule.

1 year = 5.5*10^13 kjoule.

from the data,

coal = 38 kj/g

amount of coal was burned = (5.5*10^13)/38

              = 1.45*10^12 grams